### v1

@@ 2 The basic JDSVD method
We review the basic JDSVD method [7] that is based on the standard extraction to compute the singular triplet $(\sigma_{1},u_{1},v_{1})$ of $A$.
@@
(2.1) $\qquad A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i}\perp\mathcal{U}\quad\mbox{and}\quad A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i}\perp\mathcal{V},\qquad\qquad i=1,\dots,k.$
The $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ are called the Ritz approximations of $A$ with respect to the left and right subspaces $\mathcal{U}$ and $\mathcal{V}$. In [7], condition (2.1) is denoted as $(A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i},A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i})\perp\perp(\mathcal{U},\mathcal{V}),\ i=1,\dots,k.$ Label $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. The method takes $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ as an approximation to $(\sigma_{1},u_{1},v_{1})$.
Let $\widetilde{U}\in\mathbb{R}^{M\times k}$ and $\widetilde{V}\in\mathbb{R}^{N\times k}$ be orthonormal basis matrices of $\mathcal{U}$ and $\mathcal{V}$. Then the left and right Ritz vectors $\tilde{u}_{i}=\widetilde{U}c_{i}$ and $\tilde{v}_{i}=\widetilde{V}d_{i}$ with unit-length $c_{i}\in\mathbb{R}^{k}$ and $d_{i}\in\mathbb{R}^{k}$. Define $H=\widetilde{U}^{T}A\widetilde{V}$. Then (2.1) is equivalent to
(2.2) $\qquad Hd_{i}=\theta_{i}c_{i}\qquad\mbox{and}\qquad H^{T}c_{i}=\theta_{i}d_{i},\qquad\qquad i=1,\dots,k.$
@@
(2.3) $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\theta_{1},\widetilde{U}c_{1},\widetilde{V}d_{1}),$
@@
(2.4) $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=\begin{bmatrix}A\tilde{v}_{1}-\theta_{1}\tilde{u}_{1}\\ A^{T}\tilde{u}_{1}-\theta_{1}\tilde{v}_{1}\end{bmatrix}.$
By (2.1), we particularly have
(2.5) $r\perp\perp(\tilde{u}_{1},\tilde{v}_{1}).$
@@
(2.7) $\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r$
for $(s,t)\perp\perp(\tilde{u}_{1},\tilde{v}_{1})$ approximately using the preferable MINRES method, and then uses the approximate solution $\tilde{s}$ and $\tilde{t}$ to expand $\mathcal{U}$ and $\mathcal{V}$, respectively. Specifically, we orthonormalize $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ and obtain the expansion vectors
$u_{+}=\frac{(I-\widetilde{U}\widetilde{U}^{T})\tilde{s}}{\|(I-\widetilde{U}\widetilde{U}^{T})\tilde{s}\|}\qquad\mbox{and}\qquad v_{+}=\frac{(I-\widetilde{V}\widetilde{V}^{T})\tilde{t}}{\|(I-\widetilde{V}\widetilde{V}^{T})\tilde{t}\|}.$
We then form the orthonormal basis matrices
(2.8) $\widetilde{U}_{\mathrm{new}}=[\widetilde{U},u_{+}]\qquad\mbox{and}\qquad\widetilde{V}_{\mathrm{new}}=[\widetilde{V},v_{+}],$
and generate the expanded $(k+1)$-dimensional left and right searching subspaces
$\mathcal{U}_{\mathrm{new}}=\mathrm{span}\{\widetilde{U}_{\mathrm{new}}\}=\mathrm{span}\{\widetilde{U},\tilde{s}\}\quad\mbox{and}\quad\mathcal{V}_{\mathrm{new}}=\mathrm{span}\{\widetilde{V}_{\mathrm{new}}\}=\mathrm{span}\{\widetilde{V},\tilde{t}\}.$
The basic JDSVD seeks for a new Ritz approximation to $(\sigma_{1},u_{1},v_{1})$ with respect to $\mathcal{U}_{\mathrm{new}}$ and $\mathcal{V}_{\mathrm{new}}$, and proceeds in such a way until $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ converges.
3 Convergence analysis on MINRES for solving (2.7)
Observe that the coefficient matrix of system (2.7) is the restriction of $B=\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ to the subspace
@@
(3.1) $\widetilde{W}_{1}=\begin{bmatrix}\widetilde{P}_{1}&\\ &\widetilde{Q}_{1}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2)}$
form an orthonormal basis of $(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, and the orthogonal projector in (2.7) is
@@
Denote by $w=(s^{T},t^{T})^{T}$. Then the correction equation (2.7) is
(3.2) $\widetilde{B}_{1}w=-r,$
@@
(3.3) $\widetilde{B}_{1}=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\quad\ \mbox{with}\quad\ \widetilde{B}_{1}^{\prime}=\widetilde{W}_{1}^{T}B\widetilde{W}_{1}=\begin{bmatrix}-\tau I_{M-1}&\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\\ \widetilde{Q}_{1}^{T}A^{T}\widetilde{P}_{1}&-\tau I_{N-1}\end{bmatrix}.$
Let $\tilde{w}_{j}=(\tilde{s}_{j}^{T},\tilde{t}_{j}^{T})^{T}$ be the approximate solution of (3.2) computed by MINRES at the $j$th iteration. Then the residual of $\tilde{w}_{j}$ is
(3.4) $r_{\mathrm{in},j}=\widetilde{B}_{1}\tilde{w}_{j}+r.$
We now present an equivalent form of (2.7) that will be used later.
@@
Let $\widetilde{B}_{1}^{\prime}$ be defined as in (3.3). The solution $w=\widetilde{W}_{1}z$ of (2.7), where $z\in\mathbb{R}^{M+N-2}$ is the solution to the linear system
(3.5) $\widetilde{B}_{1}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{1}^{T}r.$
If MINRES is applied to (3.2) and (3.5) with the initial guesses $\tilde{w}_{0}$ and $\tilde{z}_{0}$ related by $\widetilde{w}_{0}=\widetilde{W}_{1}\tilde{z}_{0}$, then the $j$th approximate solutions $\tilde{w}_{j}$ and $\tilde{z}_{j}$ satisfy $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$, $j=1,2,\ldots,M+N-2$. Write $r_{\mathrm{in},j}^{\prime}=\widetilde{B}_{1}^{\prime}\tilde{z}_{j}+r^{\prime}$. Then
(3.6) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|,\qquad j=0,1,\ldots,M+N-2.$
Proof.
By (2.1) and (2.7), since $r\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$ and $w\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, we have
@@
with $r^{\prime}=\widetilde{W}_{1}^{T}r$ and $z=\widetilde{W}_{1}^{T}w$. Substituting them and (3.3) into (3.2) yields
(3.7) $\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\widetilde{W}_{1}z=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}z=-\widetilde{W}_{1}r^{\prime},$
premultiplying which by $\widetilde{W}_{1}^{T}$ leads to (3.5).
@@
(3.8) $\|r_{\mathrm{in},j}\|=\|\widetilde{B}_{1}\tilde{w}_{j}+r\|=\min_{\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})}\|\widetilde{B}_{1}\tilde{w}+r\|,$
by definition it is straightforward to verify from (3.3) and (3.4) that
@@
that is, for any $\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})$, there exists $\tilde{z}\in\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime})$ such that $\tilde{w}=\widetilde{W}_{1}\tilde{z}$ and thus $\tilde{z}=\widetilde{W}_{1}^{T}\tilde{w}$, $j=1,2,\ldots,M+N-2$. Therefore, from (3.8) we have
@@
which proves (3.6) and $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$. ∎
Theorem 3.1 shows that the convergence of MINRES for (3.2) and (3.5) is identical. Without loss of generality, we take initial guesses $\tilde{w}_{0}=\bm{0}$ and $\tilde{z}_{0}=\bm{0}$ in MINRES for (3.2) and (3.5), so that their initial residuals are
(3.9) $r_{\mathrm{in},0}=r\qquad\mbox{and}\qquad r_{\mathrm{in},0}^{\prime}=r^{\prime}.$
Denote by $\tilde{\lambda}_{i},i=1,\dots,M+N-2$ the eigenvalues of $\widetilde{B}_{1}^{\prime}$ in (3.3). It is known from [5] that the residual norm of the approximate solution $\tilde{z}_{j}$ of (3.5) satisfies
(3.10) $\frac{\|r_{\mathrm{in},j}^{\prime}\|}{\|r^{\prime}\|}\leq\min_{\pi\in\Pi_{j}}\max_{1\leq i\leq M+N-2}\left|\pi\left(\tilde{\lambda}_{i}\right)\right|,$
where $\Pi_{j}$ denotes the set of polynomials of degree no more than $j$ with $\pi(0)=1$ for $\pi\in\Pi_{j}$. Keep in mind (3.6). Then there are the following estimates [5, Chapter 3].
Theorem 3.2 ([5, Chapter 3]).
@@
(3.11) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{\beta}/\sqrt{\alpha}}\right)^{j};$
@@
(3.12) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{\beta_{1}\beta_{2}}/\sqrt{\alpha_{1}\alpha_{2}}}\right)^{\left[\frac{j}{2}\right]},$
@@
This theorem indicates that the smaller $\beta-\alpha$ or $\beta_{1}-\alpha_{1}$ is, the closer is $\beta/\alpha$ or $(\beta_{1}\beta_{2})/(\alpha_{1}\alpha_{2})$ to one and thus the faster MINRES converges. Bound (3.11) is the same as that for the CG method that solves the same linear system [5], and the convergence factor is $\frac{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}-1}{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}+1}$ with $\kappa(\widetilde{B}_{1}^{\prime})=\beta/\alpha$. Suppose that $[-\beta_{1},-\alpha_{1}]$ and $[\alpha_{2},\beta_{2}]$ are symmetric with respect to the origin, i.e., $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. Then the factor in the right hand side of (3.12) is $\frac{\kappa(\widetilde{B}_{1}^{\prime})-1}{\kappa(\widetilde{B}_{1}^{\prime})+1}$, the same as that in the steepest gradient method for a positive definite (3.5) with the eigenvalue interval $[\alpha,\beta]$, but the power is $\left[\frac{j}{2}\right]$ rather than $j$. As a result, MINRES may converge much more slowly for a highly indefinite linear system than it does for a definite one.
Before proceeding, we should notice that $r=\mathbf{0}$ when $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\sigma_{1},u_{1},v_{1})$. In this case we terminate JDSVD. For later use when $\tilde{u}_{1}\not=u_{1}$ and $\tilde{v}_{1}\not=v_{1}$, we next suppose any virtual right hand side $r\not=\mathbf{0}$ satisfying (2.5). Based on Theorems 3.13.2, we present the following convergence results on MINRES for (2.7).
@@
For $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$ in (2.7), and suppose virtually $r\not=\mathbf{0}$ satisfying $r\perp\perp(u_{1},v_{1})$. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,2})/2$ with $\sigma_{\max,2}$ being the second largest singular value of $A$, then $\sigma_{1}=\sigma_{\max}$ and
(3.13) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2})/(\tau-\sigma_{\max,2})}}\right)^{j};$
(ii) if $\tau<(\sigma_{\min}+\sigma_{\min,2})/2$ with $\sigma_{\min,2}$ being the second smallest singular value of $A$, then $\sigma_{1}=\sigma_{\min}$ and
(3.14) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\left(1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,2}^{2}-\tau^{2})}}\right)^{\left[\frac{j-j_{o}}{2}\right]},$
@@
(3.15) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\tau-\sigma_{2}|}\right)^{\left[\frac{j}{2}\right]}.$
Proof.
By (1.1), for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$, take $\widetilde{P}_{1}=P_{1}$ and $\widetilde{Q}_{1}=Q_{1}$ in (3.1) with
(3.16) $P_{1}=[u_{2},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{1}=[v_{2},\dots,v_{N}],$
and rewrite the resulting $\widetilde{B}_{1}^{\prime}=B_{1}^{\prime}$ in (3.3). Then
[…]
@@
(3.18) ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
Case (i): If $\tau>(\sigma_{\max}+\sigma_{\max,2})/2$, then clearly $\sigma_{\max}$ is the closest to $\tau$, i.e., $\sigma_{1}=\sigma_{\max}$. In this case, all the singular values but $\sigma_{1}$ of $A$ are smaller than $\tau$, and
(3.19) $\lambda_{i}\in[-(\sigma_{\max,2}+\tau),-(\tau-\sigma_{\max,2})]$
by noticing that $\sigma_{2}=\sigma_{\max,2}$. Exploiting (3.11), we obtain (3.13).
Case (ii): If $\tau<(\sigma_{\min}+\sigma_{\min,2})/2$, then $\sigma_{1}=\sigma_{\min}$, and
@@
$q(\lambda)=\left(\frac{\tau+\lambda}{\tau}\right)^{j_{o}}\pi(\lambda)\qquad\mbox{with}\qquad\pi\in\Pi_{j-j_{o}},$
where $j_{o}=\min\{1,M-N\}$. Since $q(0)=\pi(0)=1$ and
$\left|\frac{\tau+\lambda}{\tau}\right|^{j_{o}}\leq\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}$
over the eigenvalues in (3.18), by Theorem 3.1 and relations (3.9)–(3.10), we obtain

### v2

@@ 2 The basic JDSVD method
We review the basic standard extraction-based JDSVD method [6] that computes the singular triplet $(\sigma_{1},u_{1},v_{1})$ of $A$.
@@
(3) $\qquad A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i}\perp\mathcal{U}\quad\mbox{and}\quad A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i}\perp\mathcal{V},\qquad i=1,\dots,k.$
The $(\theta_{i},\tilde{u}_{i},\tilde{v}_{i})$ are the Ritz approximations of $A$ with respect to $\mathcal{U}$ and $\mathcal{V}$, and (3) is denoted as $(A\tilde{v}_{i}-\theta_{i}\tilde{u}_{i},A^{T}\tilde{u}_{i}-\theta_{i}\tilde{v}_{i})\perp\perp(\mathcal{U},\mathcal{V}),\ i=1,\dots,k.$ Label $|\theta_{1}-\tau|\leq|\theta_{2}-\tau|\leq\dots\leq|\theta_{k}-\tau|$. The method takes $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})$ as an approximation to $(\sigma_{1},u_{1},v_{1})$.
Let $\widetilde{U}\in\mathbb{R}^{M\times k}$ and $\widetilde{V}\in\mathbb{R}^{N\times k}$ be orthonormal basis matrices of $\mathcal{U}$ and $\mathcal{V}$. Then the left and right Ritz vectors $\tilde{u}_{i}=\widetilde{U}c_{i}$ and $\tilde{v}_{i}=\widetilde{V}d_{i}$ with unit-length $c_{i}\in\mathbb{R}^{k}$ and $d_{i}\in\mathbb{R}^{k}$. Define $H=\widetilde{U}^{T}A\widetilde{V}$. Then (3) is equivalent to
(4) $\qquad Hd_{i}=\theta_{i}c_{i}\qquad\mbox{and}\qquad H^{T}c_{i}=\theta_{i}d_{i},\qquad i=1,\dots,k.$
@@
(5) $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\theta_{1},\widetilde{U}c_{1},\widetilde{V}d_{1}),$
@@
(6) $r=r(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=\begin{bmatrix}A\tilde{v}_{1}-\theta_{1}\tilde{u}_{1}\\ A^{T}\tilde{u}_{1}-\theta_{1}\tilde{v}_{1}\end{bmatrix}.$
By (3), we particularly have
(7) $r\perp\perp(\tilde{u}_{1},\tilde{v}_{1}).$
@@
(9) $\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}-\tau I&A\\ A^{T}&-\tau I\end{bmatrix}\begin{bmatrix}I-\tilde{u}_{1}\tilde{u}_{1}^{T}&\\ &I-\tilde{v}_{1}\tilde{v}_{1}^{T}\end{bmatrix}\begin{bmatrix}s\\ t\end{bmatrix}=-r$
for $(s,t)\perp\perp(\tilde{u}_{1},\tilde{v}_{1})$ approximately using, e.g., the preferable MINRES method, and then orthonormalizes $\tilde{s}$ and $\tilde{t}$ against $\widetilde{U}$ and $\widetilde{V}$ to obtain $u_{+}$ and $v_{+}$. The columns of
$\widetilde{U}_{\operatorname{new}}=[\widetilde{U},u_{+}]\qquad\mbox{and}\qquad\widetilde{V}_{\operatorname{new}}=[\widetilde{V},v_{+}]$
are used to construct the expanded $(k+1)$-dimensional left and right subspaces
$\mathcal{U}_{\operatorname{new}}=\operatorname{span}\{\widetilde{U}_{\operatorname{new}}\}=\operatorname{span}\{\widetilde{U},\tilde{s}\}\quad\mbox{and}\quad\mathcal{V}_{\operatorname{new}}=\operatorname{span}\{\widetilde{V}_{\operatorname{new}}\}=\operatorname{span}\{\widetilde{V},\tilde{t}\}.$
JDSVD then computes a new approximation to $(\sigma_{1},u_{1},v_{1})$ with respect to $\mathcal{U}_{\operatorname{new}}$ and $\mathcal{V}_{\operatorname{new}}$.
3 Convergence analysis on MINRES for solving (9)
Observe that the coefficient matrix of system (9) is the restriction of $B=\begin{bmatrix}\begin{smallmatrix}-\tau I&A\\ A^{T}&-\tau I\end{smallmatrix}\end{bmatrix}$ to the subspace
@@ 3 Convergence analysis on MINRES for solving (9)
(10) $\widetilde{W}_{1}=\begin{bmatrix}\widetilde{P}_{1}&\\ &\widetilde{Q}_{1}\end{bmatrix}\in\mathbb{R}^{(M+N)\times(M+N-2)}$
form an orthonormal basis of $(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, and the orthogonal projector in (9) is
@@
Denote by $w=(s^{T},t^{T})^{T}$. Then the correction equation (9) is
(11) $\widetilde{B}_{1}w=-r,$
@@
(12) $\widetilde{B}_{1}=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\quad\ \mbox{with}\quad\ \widetilde{B}_{1}^{\prime}=\widetilde{W}_{1}^{T}B\widetilde{W}_{1}=\begin{bmatrix}-\tau I_{M-1}&\widetilde{P}_{1}^{T}A\widetilde{Q}_{1}\\ \widetilde{Q}_{1}^{T}A^{T}\widetilde{P}_{1}&-\tau I_{N-1}\end{bmatrix}.$
Let $\tilde{w}_{j}=(\tilde{s}_{j}^{T},\tilde{t}_{j}^{T})^{T}$ be the approximate solution of (11) computed by MINRES at the $j$th iteration. Then the residual of $\tilde{w}_{j}$ is
(13) $r_{\mathrm{in},j}=\widetilde{B}_{1}\tilde{w}_{j}+r.$
We now present an equivalent form of (9) that will be used later.
@@
Let $\widetilde{B}_{1}^{\prime}$ be defined as in (12). Then the solution $w=\widetilde{W}_{1}z$ to equation (9), where $z\in\mathbb{R}^{M+N-2}$ is the solution to the linear system
(14) $\widetilde{B}_{1}^{\prime}z=-r^{\prime}\qquad\mbox{with}\qquad r^{\prime}=\widetilde{W}_{1}^{T}r.$
If MINRES is applied to (11) and (14) with the initial guesses $\tilde{w}_{0}$ and $\tilde{z}_{0}$ related by $\widetilde{w}_{0}=\widetilde{W}_{1}\tilde{z}_{0}$, the $j$th approximate solutions $\tilde{w}_{j}$ and $\tilde{z}_{j}$ obtained by it satisfy $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$, $j=1,2,\ldots,M+N-2$. Write $r_{\mathrm{in},j}^{\prime}=\widetilde{B}_{1}^{\prime}\tilde{z}_{j}+r^{\prime}$. Then the residual norms
(15) $\|r_{\mathrm{in},j}\|=\|r^{\prime}_{\mathrm{in},j}\|,\qquad j=0,1,\ldots,M+N-2.$
Proof 3.2.
By (7) and (9), since $r\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$ and $w\in(\tilde{u}_{1},\tilde{v}_{1})^{\perp\perp}$, we have
@@
with $r^{\prime}=\widetilde{W}_{1}^{T}r$ and $z=\widetilde{W}_{1}^{T}w$. Substituting them and (12) into (11) yields
(16) $\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}\widetilde{W}_{1}^{T}\widetilde{W}_{1}z=\widetilde{W}_{1}\widetilde{B}_{1}^{\prime}z=-\widetilde{W}_{1}r^{\prime},$
premultiplying which by $\widetilde{W}_{1}^{T}$ leads to (14).
@@
(17) $\|r_{\mathrm{in},j}\|=\|\widetilde{B}_{1}\tilde{w}_{j}+r\|=\min_{\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})}\|\widetilde{B}_{1}\tilde{w}+r\|,$
by definition it is straightforward to verify from (12) and (13) that
@@
that is, for any $\tilde{w}\in\mathcal{K}_{j}(\widetilde{B}_{1},r_{\mathrm{in},0})$, there exists $\tilde{z}\in\mathcal{K}_{j}(\widetilde{B}_{1}^{\prime},r_{\mathrm{in},0}^{\prime})$ such that $\tilde{w}=\widetilde{W}_{1}\tilde{z}$ and thus $\tilde{z}=\widetilde{W}_{1}^{T}\tilde{w}$, $j=1,2,\ldots,M+N-2$. Therefore, from (17) we obtain
@@
which proves (15) and $\tilde{w}_{j}=\widetilde{W}_{1}\tilde{z}_{j}$.
Theorem 3.1 shows that the convergence of MINRES for (11) and (14) is identical. Without loss of generality, take $\tilde{w}_{0}=\bm{0}$ and $\tilde{z}_{0}=\bm{0}$. Then
(18) $r_{\mathrm{in},0}=r\qquad\mbox{and}\qquad r_{\mathrm{in},0}^{\prime}=r^{\prime}.$
Denote by $\tilde{\lambda}_{i},i=1,\dots,M+N-2$ the eigenvalues of $\widetilde{B}_{1}^{\prime}$ in (12). It is known from [4] that the residual norm of the approximate solution $\tilde{z}_{j}$ of (14) satisfies
(19) $\frac{\|r_{\mathrm{in},j}^{\prime}\|}{\|r^{\prime}\|}\leq\min_{\pi\in\Pi_{j}}\max_{1\leq i\leq M+N-2}|\pi(\tilde{\lambda}_{i})|,$
where $\Pi_{j}$ denotes the set of polynomials of degree no more than $j$ and $\pi(0)=1$ for $\pi\in\Pi_{j}$. Keep in mind (15). Then there are the following estimates.
Theorem 3.3 ([4, Chapter 3]).
@@
(20) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\tfrac{2}{1+\sqrt{\beta}/\sqrt{\alpha}}\right)^{j};$
@@
(21) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{\beta_{1}\beta_{2}}/\sqrt{\alpha_{1}\alpha_{2}}}\right)^{\left[\frac{j}{2}\right]},$
@@
This theorem indicates that the smaller $\beta-\alpha$ or $\beta_{1}-\alpha_{1}$ is, the closer $\beta/\alpha$ or $(\beta_{1}\beta_{2})/(\alpha_{1}\alpha_{2})$ is to one and thus the faster MINRES converges. Bound (20) is the same as that for the CG method for solving the same linear system [4], where the convergence factor is $\frac{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}-1}{\sqrt{\kappa(\widetilde{B}_{1}^{\prime})}+1}$ with $\kappa(\widetilde{B}_{1}^{\prime})=\beta/\alpha$. Suppose that $[-\beta_{1},-\alpha_{1}]$ and $[\alpha_{2},\beta_{2}]$ are symmetric with respect to the origin, i.e., $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. Then the factor in the right-hand side of (21) is $\frac{\kappa(\widetilde{B}_{1}^{\prime})-1}{\kappa(\widetilde{B}_{1}^{\prime})+1}$, the same as that for the steepest gradient method applied to a positive definite (14) with the eigenvalue interval $[\alpha,\beta]$, but the power is $\left[\frac{j}{2}\right]$ rather than $j$. As a result, MINRES may converge much more slowly for a highly indefinite linear system than it does for a definite one.
Before proceeding, we should notice that $r=\mathbf{0}$ when $(\theta_{1},\tilde{u}_{1},\tilde{v}_{1})=(\sigma_{1},u_{1},v_{1})$. In this case we terminate JDSVD. For later use, we next suppose an arbitrary virtual right-hand side $r\not=\mathbf{0}$ satisfying (7) when $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$. Based on Theorems 3.13.3, we present the following convergence results on MINRES for (9).
@@
For $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$ in (9), and suppose virtually $r\not=\mathbf{0}$ satisfying $r\perp\perp(u_{1},v_{1})$. Then (i) if $\tau>(\sigma_{\max}+\sigma_{\max,2})/2$ with $\sigma_{\max,2}$ being the second largest singular value of $A$, then $\sigma_{1}=\sigma_{\max}$ and
(22) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+\sqrt{(\tau+\sigma_{\max,2})/(\tau-\sigma_{\max,2})}}\right)^{j};$
(ii) if $\sigma_{\min}\leq\tau<(\sigma_{\min}+\sigma_{\min,2})/2$ with $\sigma_{\min,2}$ being the second smallest singular value of $A$, then $\sigma_{1}=\sigma_{\min}$ and
(23) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}\Bigg{(}1-\frac{2}{1+\sqrt{(\sigma_{\max}^{2}-\tau^{2})/(\sigma_{\min,2}^{2}-\tau^{2})}}\Bigg{)}^{\left[\frac{j-j_{o}}{2}\right]},$
@@
(24) $\frac{\|r_{\mathrm{in},j}\|}{\|r\|}\leq 2\left(1-\frac{2}{1+(\sigma_{\max}+\tau)/|\tau-\sigma_{2}|}\right)^{\left[\frac{j}{2}\right]}.$
Proof 3.5.
By (1), for $\tilde{u}_{1}=u_{1}$ and $\tilde{v}_{1}=v_{1}$, take $\widetilde{P}_{1}=P_{1}$ and $\widetilde{Q}_{1}=Q_{1}$ in (10) with
(25) $P_{1}=[u_{2},\dots,u_{M}]\qquad\mbox{and}\qquad Q_{1}=[v_{2},\dots,v_{N}],$
and rewrite the resulting $\widetilde{B}_{1}^{\prime}=B_{1}^{\prime}$ in (12). Then
(26) $B_{1}^{\prime}=\begin{bmatrix}-\tau I_{M-1}&\Sigma_{1}^{\prime}\\ {\Sigma_{1}^{\prime}}^{T}&-\tau I_{N-1}\end{bmatrix}\qquad\mbox{with}\qquad\Sigma_{1}^{\prime}=P_{1}^{T}AQ_{1}=\operatorname{diag}\{\sigma_{2},\dots,\sigma_{N}\}.$
@@
(27) ±$\pm\sigma_{2}-\tau,\dots,\pm\sigma_{N}-\tau\qquad\mbox{and}\qquad\underbrace{-\tau,\dots,-\tau}_{M-N}.$
Case (i): In this case, $\sigma_{\max}$ is the closest to $\tau$, i.e., $\sigma_{1}=\sigma_{\max}$, all the singular values but $\sigma_{1}$ of $A$ are smaller than $\tau$, and
(28) $\lambda_{i}\in[-(\sigma_{\max,2}+\tau),-(\tau-\sigma_{\max,2})]$
by noticing that $\sigma_{2}=\sigma_{\max,2}$. Exploiting (20), we obtain (22).
Case (ii): In this case, we have $\sigma_{1}=\sigma_{\min}$ and
@@
$q(\lambda)=\left(\tfrac{\tau+\lambda}{\tau}\right)^{j_{o}}\pi(\lambda)\qquad\mbox{with}\qquad\pi\in\Pi_{j-j_{o}},$
where $j_{o}=\min\{1,M-N\}$. Since $q(0)=\pi(0)=1$ and $\left|\frac{\tau+\lambda}{\tau}\right|^{j_{o}}\leq\left(\frac{\sigma_{\max}}{\tau}\right)^{j_{o}}$ over the eigenvalues in (27), by Theorem 3.1 and relations (18)–(19), we obtain